OpenStudy (anonymous):

what is the [H+] of a solution with a pH of 6.6? Use a Graphic calculator

OpenStudy (janu16):

got options?

OpenStudy (anonymous):

2.5 × 107 5.0 × 105 8.3 × 106 7.7 × 105

OpenStudy (janu16):

[H+] = 10-pH = 10-6.6 which will be between 1.0x10-7M and 1.0x10-6M which still leaves us with 2 possible answer choices (2.5x10-7M and 6.6x10-7M) so we’ll have to be more precise. You’ll have to remember not only the [H+] values when the pH is an integer but also for the halves (.5) too. The half pH occurs when [H+] = 3.16x10-whatever as demonstrated below: [H+] = 1x10-6M pH = 6 [H+] = 3.16x10-7M pH = 6.5 [H+] = 1x10-7M pH = 7 Since a pH of 6.6 is between 6.5 and 7, the [H+] must be between 1x10-7M and 3.16x10-7M and therefore 2.5x10-7M must be the correct answer.

OpenStudy (janu16):

so 2.5 x 10^-7

OpenStudy (anonymous):

oh ok thank you !!